OpenStudy (anonymous):

Write the equation of the line that goes through the point (-1, 4) and is perpendicular to the line 2x + 3y – 8 = 0 in general form.

OpenStudy (jdoe0001):

what's the slope of => 2x + 3y – 8 = 0 ?

OpenStudy (jdoe0001):

well, solve the equation first for "y" see what you get

OpenStudy (anonymous):

y = (-2/3)x + (8/3) this should be the answer for Y x = 4

OpenStudy (jdoe0001):

well, no quite yet that's just to get the slope for that equation \(\begin{matrix} y=&\color{red}{-\cfrac{2}{3}}x&+\cfrac{8}{3}\\ & \color{red}{slope}& \end{matrix}\)

OpenStudy (jdoe0001):

so the slope of that equation is -2/3 the slope for a PERPENDICULAR line to that one, will be NEGATIVE RECIPROCAL of that one what does that mean? this one is -2/3 RECIPROCAL of that -3/2 NEGATIVIZE it 3/2

OpenStudy (anonymous):

So that would be the answer? Or just the slope?

OpenStudy (jdoe0001):

so now we know that, the other line has a slope of 3/2 and goes through the point (-1, 4) so we use that and plug it in the point-slope form of \(\bf y-y_1=m(x-x_1)\) \(\bf (x_1, y_1) \implies (-1, 4)\)

OpenStudy (jdoe0001):

m = slope

OpenStudy (anonymous):

M= 3/2

OpenStudy (jdoe0001):

\(\bf m = slope = \cfrac{3}{2}\\ (x_1, y_1) = (-1,4)\\ \color{blue}{y-y_1=m(x-x_1)}\implies y-(4)=\cfrac{3}{2}(x-(-1))\) then again, just solve for "y", to get the equation :)

OpenStudy (anonymous):

Y - (4) = 3/2 (x -(-1))? and find the value of Y?

OpenStudy (jdoe0001):

yes, so-called finding or solving for "y", is really just going to make it in y = ax+b form which is the so-called slope-intercept form of the equation, which I think it's what's expected

OpenStudy (anonymous):

How do i find Y?

OpenStudy (jdoe0001):

\(\bf y-(4)=\cfrac{3}{2}(x-(-1)) \implies y-4 = \cfrac{3(x+1)}{2}\\ 2y-8 = 3(x+1) \implies \square? = \square?\)

OpenStudy (jdoe0001):

\(\bf 2y-8 = 3(x+1) \implies 2y-8 = 3x+3 \implies \square? = \square?\)